# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

### v3

@@ 3.2 Dataset Details
Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 / 46.30 73.00 79.15 88.60 83.83 85.15
@@ 4.2 Main Results
Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the Appendix. As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks.
@@
In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86), ML (81.90) and DC (82.78), showing superior multi-disciplinary knowledge and reasoning. MiMo-Audio performs strongly in ML (81.75), comparable to Qwen3-Omni. However, Baichuan-Audio (GK:40.73, ML:74.05), GLM-4-Voice (GK:41.79, ML:60.18) and other models lag, revealing gaps in knowledge coverage and step-by-step reasoning.
Overall, Qwen3-Omni demonstrates comprehensive superiority. Other models have task-specific strengths: Fun-Audio-Chat in cross-lingual TIF and SIF, MiMo-Audio in TIF and ML. Most open-source E2E LALMs face English adaptation challenges, while GPT-4o-Audio lacks long-context dialogue capability.
@@ A.3 Complete Results Of Knowledge
For Mathematical and Logical Reasoning, Qwen3-Omni (81.90) and MiMo-Audio (81.75) stand out with the highest averages. Qwen3-Omni dominates in Logic (95.60) and Induction (90.62), while MiMo-Audio tops Medium Math (87.06). However, Models like GLM-4-Voice (60.18) and Fun-Audio-Chat (73.00) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
For Discourse Comprehension, Qwen3-Omni (82.78) achieves the highest average, excelling in Induction (92.04), Inference (89.32) and Analysis (67.83). Step-Audio 2 mini (79.46) and Fun-Audio-Chat (79.15) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
Overall, Qwen3-Omni demonstrates robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning. Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.

### v4

@@ 3.2 Dataset Details
Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 85.27 53.89 86.12 87.31 88.60 83.83 85.15
@@ 4.2 Main Results
Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the Appendix. As shown, Qwen3-Omni and Fun-Audio-Chat emerge as the new state-of-the-art (SOTA) models, demonstrating all-round superiority across most tasks.
@@
In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86) and Fun-Audio-Chat performs strongly in ML (86.12) and DC (87.31), showing superior multi-disciplinary knowledge and reasoning. However, Baichuan-Audio (GK:40.73, ML:74.05), GLM-4-Voice (GK:41.79, ML:60.18) and other models lag, revealing gaps in knowledge coverage and step-by-step reasoning.
Overall, Qwen3-Omni and Fun-Audio-Chat demonstrate comprehensive superiority, while other models have task-specific strengths. Most open-source E2E LALMs face English adaptation challenges, while GPT-4o-Audio lacks long-context dialogue capability.
@@ A.3 Complete Results Of Knowledge
For Mathematical and Logical Reasoning, Fun-Audio-Chat (86.12) and Qwen3-Omni (81.90) stand out with the highest averages. Qwen3-Omni dominates in logical reasoning, while MiMo-Audio tops math. However, models like GLM-4-Voice (60.18) shows weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
For Discourse Comprehension, Fun-Audio-Chat (87.31) achieves the highest average, excelling in Induction (88.50), Inference (95.15) and Analysis (79.13). Qwen3-Omni (82.78) and Step-Audio 2 mini (79.46) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
Overall, Qwen3-Omni and Fun-Audio-Chat demonstrate robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning. Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.